# Phantom: General Trigger Attacks on Retrieval Augmented Language Generation

### v2

@@ 4 Phantom Attack Framework
We now construct our Phantom attack framework to induce the adversarial objectives outlined in Section 3.1. In Phantom, the adversary executes a two-fold attack: i) poisoning the RAG retriever followed by ii) compromising the RAG generator. To achieve this dual objective the adversary creates an adversarial passage $\mathsf{p_{adv}}=\mathsf{s_{ret}}\oplus\mathsf{s_{gen}}\oplus\mathsf{s_{cmd}}$, where $\oplus$ denotes string concatenation, and inserts it into a single document in the victim’s local knowledge base. We ensure that the size of the adversarial passage is below of the passage length used by the retriever, so that the entire adversarial passage can be retrieved at once. Here, $\mathsf{s_{ret}}$ represents the adversarial payload for the retriever component, while $\mathsf{s_{gen}}$ targets the generator and $\mathsf{s_{cmd}}$ is the adversarial command.
@@
When attacking the retriever, the adversary’s aim is to ensure that the adversarial passage is chosen among the top-$k$ documents selected by the retriever, but only when the trigger sequence $\mathsf{s_{trg}}$ appears in the user’s query. For this purpose, we propose a new optimization approach that constructs the adversarial retriever string $\mathsf{s_{ret}}$ to maximize the likelihood of $\mathsf{p_{adv}}$ being in the top-$k$ passages.
@@
The next step involves creating the adversarial generator string $\mathsf{s_{gen}}$, used to break the alignment of the LLM generator. The goal is to execute the adversarial command $\mathsf{s_{cmd}}$ and output a response starting with a target string $\mathsf{s_{op}}$. We propose an extension to GCG 59 to break the LLM alignment with faster convergence. Figure 2 provides a visual illustration our framework pipeline.
4.1 Attacking the RAG Retriever
@@
The adversary aims to construct a retriever string $\mathsf{s_{ret}}$ to ensure that adversarial passage $\mathsf{p_{adv}}$ is selected in the top-$k$ passages if and only if the specified trigger sequence $\mathsf{s_{trg}}$ is present within a user’s query $\mathsf{q^{in}_{j}}$. One approach for creating $\mathsf{s_{ret}}$ could be to just repeat the trigger $\mathsf{s_{trg}}$ multiple times. However, this simple baseline fails, as shown in Section A.3.1, because the adversarial passage never appears in the top-$k$ retrieved documents, highlighting the need for an optimization procedure. In Section A.4, we also explore how to predict whether a trigger itself will be viable for use in our attack.
@@
i) Loss Modelling. We start with creating an OUT set $\{\mathsf{q^{out}_{1}},\ldots,\mathsf{q^{out}_{n}}\}$ composed of queries without the trigger sequence $\mathsf{s_{trg}}$. Note that these queries are not necessarily related to each other. Next, we append $\mathsf{s_{trg}}$ to each query and create our IN set $\{\mathsf{q^{out}_{1}}\oplus\mathsf{s_{trg}},\ldots,\mathsf{q^{out}_{n}}\oplus\mathsf{s_{trg}}\}$. The loss function can then be written as:
$\mathsf{L_{ret}}=\frac{1}{\mathsf{n}}\sum_{i=1}^{\mathsf{n}}\mathsf{sim}({\mathsf{E}_{\mathsf{p_{adv}^{*}}}},{\mathsf{E}_{\mathsf{q^{out}_{j}}}})-\mathsf{sim}({\mathsf{E}_{\mathsf{p_{adv}^{*}}}},{\mathsf{E}_{\mathsf{q^{out}_{j}}\oplus\mathsf{s_{trg}}}})$ (1)
@@
We can now formulate Equation 1 as an optimization problem and solve it using the HotFlip technique 10 as follows.
ii) Optimization Algorithm. Let $\mathsf{t_{1:r}}=\{t_{1},..,t_{r}\}$ denote the tokenized version of the adversarial string $\mathsf{s_{ret}}$. HotFlip identifies candidate replacements for each token $t_{i}\in\mathsf{t_{1:r}}$ that reduces the loss $\mathsf{L_{ret}}$. The optimization starts by initializing $\mathsf{t_{1:r}}$ with mask tokens IDs, provided by the tokenizer, and at each step a token position $i$, chosen sequentially, is swapped with the best candidate computed as:
@@
where $\mathcal{V}$ denotes the vocabulary and $e_{t_{i}}$ and $\nabla_{e_{t_{i}}}$ denote the token embedding and the gradient vector with respect to $t_{i}$’s token embedding, respectively. Multiple epochs of the process can be used to iteratively update the entire $\mathsf{s_{ret}}$.
4.2 Attacking the RAG Generator
@@
Once $\mathsf{s_{ret}}$ is ready, the next step involves creating the adversarial generator string $\mathsf{s_{gen}}$. This string $\mathsf{p_{adv}}$ induces the generator to execute the adversarial command $\mathsf{s_{cmd}}$. Note that we want this to occur for any query with the trigger sequence $\mathsf{s_{trg}}$, regardless of the rest of the query’s content. To facilitate this functionality, we extend the GCG attack by 59.
@@
i) Loss Modelling. Let $\mathrm{Q_{in}}=\{\mathsf{q^{in}_{1}},\ldots,\mathsf{q^{in}_{m}}\}$ denote a set of queries with trigger sequence $\mathsf{s_{trg}}$ present in each query. We define $\mathrm{P_{top}}$ as the set of top-$k$ passages retrieved in response to a user query $\mathsf{q^{in}_{j}}\in\mathrm{Q_{in}}$, with $\mathsf{p_{adv}}$ included within $\mathrm{P_{top}}$. Let $\mathrm{T_{in}}=\{\mathsf{t_{1:w_{1}}},\ldots,\mathsf{t_{1:w_{m}}}\}$ denote the tokenized input, where $\mathsf{t_{1:w_{j}}}$ represents the tokenized version of the user query $\mathsf{q^{in}_{j}}$ and its passages $\mathrm{P_{top}}$ structured in the pre-specified RAG template format. We use $\mathsf{t_{gen}}\subset\mathsf{t_{1:w_{j}}}$ to denote the subset of tokens that correspond to the string $\mathsf{s_{gen}}$ and $\mathsf{t_{w_{j}+1:w_{j}+c}}$ to represent the tokenized version of the target string $\mathsf{s_{op}}$. We can now formulate the loss function as
$\mathsf{L_{gen}}(\mathrm{T_{in}})=-\sum_{j=1}^{\mathsf{m}}\log\Pr(\mathsf{t_{w_{j}+1:w_{j}+c}}|\mathsf{t_{1:w_{j}}})=-\sum_{j=1}^{\mathsf{m}}\sum_{i=1}^{c}\log\Pr(\mathsf{t_{w_{j}+i}}|\mathsf{t_{1:w_{j}+i-1}})$ (2)
@@
where $\Pr(\mathsf{t_{w_{j}+i}}|\mathsf{t_{1:w_{j}+i-1}})$ denotes the probability of generating token $\mathsf{t_{w_{j}+i}}{}$ given the sequence of tokens up to position $\mathsf{w_{j}+i-1}$. Our goal is to minimize the $\mathsf{L_{gen}}$ loss by optimizing the list of tokens $\mathsf{t_{gen}}$ associated to adversarial generator string $\mathsf{s_{gen}}$. We now formulate an optimization problem to solve Equation 2.
@@
ii) Optimization Algorithm. One potential strategy could be to implement the GCG optimization 59 to minimize our loss. However, we observe that GCG attack requires a large batch size and many iterations to lower the loss and break the generator’s alignment. Instead, we propose a more efficient approach, called Multi Coordinate Gradient (MCG), that reduces the number of iterations and batch size. Algorithm 1 presents out method, which simultaneously updates a subset of $C$ coordinates per iteration. We generate a batch of $B$ candidates, and for each of them we update $C$ random coordinates with one of the top-$k$ token replacements. The candidate with the lowest $\mathsf{L_{gen}}$ is then chosen as the best substitution for $\mathsf{t_{gen}}$.
Algorithm 1 Multi Coordinate Gradient (MCG)
1:Input: tokenized set $\mathrm{T_{in}}$, adversarial generator tokens $\mathsf{t_{gen}}$, total iterations $I$, generator loss $\mathsf{L_{gen}}$, number of coordinates $C$, batch size $B$, minimum number of coordinates to change per iteration $\mathsf{c_{min}}$
2:for $I$ iterations do
3:    for $t_{i}\in\mathsf{t_{gen}}$ do
4:        $S_{i}=\text{top-k}(-\nabla_{e_{\mathsf{t_{i}}}}\mathsf{L_{gen}}(\mathrm{T_{in}}))$triangleright Compute top-k token substitutions
5:    for b = 1,…,B do
6:        $\mathsf{t_{sub}^{b}}=\mathsf{t_{gen}}$
7:        for C iterations do $\triangleright$ Select random subset of C coordinates
8:           $\mathsf{t_{c}^{b}}:=\text{Uniform}(S_{c})$, where $c=\text{Uniform}(\text{len}(\mathsf{t_{sub}^{b}}))$triangleright Select token replacement
9:    $\mathsf{t_{gen}}:=\mathsf{t_{sub}^{b^{*}}}$, where $b^{*}=\operatorname*{arg\,min}_{b}\mathsf{L_{gen}}(\mathrm{T_{in}})$triangleright Select best substitution
10:    $C:=\max(C/2,\mathsf{c_{min}})$triangleright Reduce number of coordinates to replace return $\mathsf{t_{gen}}$
@@
We observe that loss fluctuates across iterations when C is large. Therefore, after each iteration, we progressively halve the number of coordinates to steadily decrease the $\mathsf{L_{gen}}$. While a large batch size $B$ (e.g., 512) and many iterations (e.g., 128) can also help decrease the loss, this approach is slower and conflicts with our efficiency goal. For success in fewer iterations (e.g., 4 to 16), changing at least 2 coordinates per iteration ($\mathsf{c_{min}}=2$ in Step 10) helps mitigate loss fluctuations while maintaining a manageable batch size of 64. Note that the GCG attack is a special case of our MCG approach when the number of coordinates is set to $C=1$. In Table 7, Section A.2, we empirically show that our approach achieves higher attack success than GCG, especially when the adversary has a small iteration budget of $I=4$ and batch size of $B=64$.
4.3 Phantom Attack Instances
@@
Table 1: Concrete Instantiation of Phantom Framework: Examples of adversarial strings $\mathsf{s_{cmd}}$ and $\mathsf{s_{op}}$ used to achieve different adversarial objectives.
@@ 5 Evaluation
All experiments are repeated three times with different random seeds, and all reported values are the averages of these results. Experiments were run on machines with at least 256GB of memory available and different GPU configurations: Nvidia RTX 4090, Nvidia RTX A6000 and Nvidia A100. Each attack can be run using a single GPU.
Datasets. To evaluate our process, we use the MS MARCO question and answer dataset 33, which contains more than 8 million document passages and roughly 1 million real user queries. For completeness, we also show our effectiveness of our attack on two other datasets, namely Natural Question (NQ) and HotPot-QA. We show these results in Section A.5 due to space constraints.
Retrievers and Generators. We evaluate our attack on a variety of commonly used, open source, retriever and generator models. In particular, we use Contriever 12, Contriever-MS (a version of Contriever fine-tuned on the MS MARCO dataset), and DPR 24 as retrievers. For the generator, we test our attack on the following LLMs: Vicuña 5 version 1.5 7B and 13B, Llama 3Currently the best ¡10B parameters open chatbot model https://chat.lmsys.org/?leaderboard Instruct 8B 49; 29, and Gemma Instruct 2B and 7B 48. Section A.2.9 presents experiments on larger closed source models.
@@
RAG Hyperparameters. For our main experiment we set the number of passages retrieved to a default value $k=5$. To evaluate the success of our attack, we conduct tests using 25 queries selected from the test set, which ensures that the trigger appears naturally within each query. Section A.2.5 and Section A.2.8 present ablation studies where we vary these parameters. We test for three triggers, for our adversarial objective, where we choose common brand or celebrity names as they frequently appear in natural user queries, and the success of our attack can cause significant reputation damage to these entities.
5.1 Phantom attack on RAG System
Here we evaluate the performance of Phantom on our adversarial objectives. Due to space constraints we defer to Appendix A the assessment of how different RAG generator and retriever parameters affect the attack’s performance. In particular, for the generator, in Section A.2, we show the impact on attack success of the MCG optimization, the number of MCG iterations and tokens and the transfer rate of our attack across generators. On the retriever side, Section A.3 examines how varying the number of HotFlip epochs, the number of retriever tokens, and different retriever architectures affect the Retrieval Failure Rate (Ret-FR).
@@ 5.1 Phantom attack on RAG System
Table 2: Effectiveness of Phantom Attack for Refusal to Answer and Biased Opinion objectives: Attack success reported for two objectives, over three triggers, in scenarios where the adversarial document is retrieved versus No-attack conditions. The RAG uses Contriever for retrieval and one of four LLMs as the generator. Symbol denotes the RAG breaks alignment with only the adversarial command $\mathsf{s_{cmd}}$, while indicates also the need of MCG optimization to break RAG’s alignment.
@@
Table 2 reports the results for our Refusal to Answer and Biased Opinion objectives. We observe that for the Refusal to Answer objective all four generators execute the command with high success rate, without requiring to run the MCG optimization. We believe the adversary’s Refusal to Answer objective aligns with RAG’s tendency to refrain from answering when uncertain about the user’s prompt, making it easier to succeed.
@@
For the Biased Opinion objective, we are able to induce Gemma-2B and Llama3-8B to produce a biased opinion, with only the adversarial command $\mathsf{s_{cmd}}$ and no MCG optimization. However, for Vicuna-7B and Gemma-7B, the adversarial command $\mathsf{s_{cmd}}$ alone is not sufficient to consistently break the alignment. Consequently, we run 16 iterations of MCG optimization over 16 tokens to successfully break the alignment of these models as well. We observe a significant improvement of $38.7\%$ and $38.4\%$ in the attack’s success across the three triggers when we append MCG’s $\mathsf{s_{gen}}$ to the adversarial command $\mathsf{s_{cmd}}$ for Vicuna-7B and Gemma-7B, respectively. A detailed comparison can be found in Table 4 (Section A.2), which shows the Biased Opinion attack success with and without MCG optimization. We include an alternative version of Table 2 in Section A.6, showing the average number of queries for which the attack was successful (with related standard deviations) across three runs. Concrete examples for our Biased Opinion objective, categorized into three groups, can be found in Section A.7.1.

### v3

@@ 4. Phantom Attack Framework
We now construct our Phantom attack framework to induce the adversarial objectives outlined in Section 3.1. In Phantom, the adversary executes a two-fold attack: i) poisoning the RAG retriever followed by ii) compromising the RAG generator. To achieve this dual objective the adversary creates an adversarial passage $\mathsf{p_{adv}}=\mathsf{s_{ret}}\oplus\mathsf{s_{gen}}\oplus\mathsf{s_{cmd}}$, where $\oplus$ denotes string concatenation, and inserts it into a single document in the victim’s local knowledge base. We ensure the size of the adversarial passage is below the passage length used by the retriever, so that the entire adversarial passage can be retrieved at once. Here, $\mathsf{s_{ret}}$ represents the adversarial payload for the retriever component, while $\mathsf{s_{gen}}$ targets the generator and $\mathsf{s_{cmd}}$ is the adversarial command.
@@
When attacking the retriever, the adversary’s aim is to ensure that the adversarial passage is chosen among the top most relevant documents selected by the retriever, but only when the trigger sequence $\mathsf{s_{trg}}$ appears in the user’s query. For this purpose, we propose a new optimization approach that constructs the adversarial retriever string $\mathsf{s_{ret}}$ to maximize the likelihood of $\mathsf{p_{adv}}$ being in the top retrieved passages.
@@
The next step involves creating the adversarial generator string $\mathsf{s_{gen}}$, used to break the alignment of the LLM generator. The goal is to execute the adversarial command $\mathsf{s_{cmd}}$ and output a response starting with a target string $\mathsf{s_{op}}$. We propose an extension to GCG (Zou et al., 2023) to break the LLM alignment with faster convergence. Figure 2 provides a visual illustration of our framework pipeline.
4.1. Attacking the RAG Retriever
@@
The adversary aims to construct a retriever string $\mathsf{s_{ret}}$ to ensure that adversarial passage $\mathsf{p_{adv}}$ is selected in the top passages if and only if the specified trigger sequence $\mathsf{s_{trg}}$ is present within a user’s query $\mathsf{q^{in}_{j}}$. One approach for creating $\mathsf{s_{ret}}$ could be to just repeat the trigger $\mathsf{s_{trg}}$ multiple times. However, this simple baseline fails (as shown in Section 6.2.1) because the adversarial passage never appears in the top-$k$ retrieved documents, highlighting the need for an optimization procedure. In Section 6.3, we also explore how to predict whether a trigger itself will be viable for use in our attack.
@@
i) Loss Modeling.
We start with creating an OUT set $\{\mathsf{q^{out}_{1}},\ldots,\mathsf{q^{out}_{n}}\}$ composed of queries without the trigger sequence $\mathsf{s_{trg}}$. Note that these queries are not necessarily related to each other. Next, we append $\mathsf{s_{trg}}$ to each query and create our IN set $\{\mathsf{q^{out}_{1}}\oplus\mathsf{s_{trg}},\ldots,\mathsf{q^{out}_{n}}\oplus\mathsf{s_{trg}}\}$. The loss function can then be written as:
(1) $\mathsf{L_{ret}}=\frac{1}{\mathsf{n}}\sum_{j=1}^{\mathsf{n}}\mathsf{sim}({\mathsf{E}_{\mathsf{p_{adv}^{*}}}},{\mathsf{E}_{\mathsf{q^{out}_{j}}}})-\mathsf{sim}({\mathsf{E}_{\mathsf{p_{adv}^{*}}}},{\mathsf{E}_{\mathsf{q^{out}_{j}}\oplus\mathsf{s_{trg}}}})$
@@
We can now formulate Equation 1 as an optimization problem and solve it using the HotFlip technique (Ebrahimi et al., 2018) as follows.
ii) Optimization Algorithm.
Let $\mathsf{t_{1:r}}=\{t_{1},..,t_{r}\}$ denote the tokenized version of the adversarial string $\mathsf{s_{ret}}$. HotFlip identifies candidate replacements for each token $t_{i}\in\mathsf{t_{1:r}}$ that reduces the loss $\mathsf{L_{ret}}$. The optimization starts by initializing $\mathsf{t_{1:r}}$ with mask tokens IDs, provided by the tokenizer, and at each step a token position $i$, chosen sequentially, is swapped with the best candidate computed as:
@@
where $\mathcal{V}$ denotes the vocabulary and $e_{t_{i}}$ and $\nabla_{e_{t_{i}}}$ denote the token embedding and the gradient vector with respect to $t_{i}$’s token embedding, respectively. Multiple epochs of the process can be used to iteratively update the entire string $\mathsf{s_{ret}}$.
4.2. Attacking the RAG Generator
@@
Once $\mathsf{s_{ret}}$ is ready, the next step involves creating the adversarial generator string $\mathsf{s_{gen}}$. This string $\mathsf{p_{adv}}$ induces the generator to execute the adversarial command $\mathsf{s_{cmd}}$. Note that we want this to occur for any query with the trigger sequence $\mathsf{s_{trg}}$, regardless of the rest of the query’s content. To facilitate this functionality, we extend the GCG attack by (Zou et al., 2023).
i) Loss Modelling.
@@
Let $\mathrm{Q_{in}}=\{\mathsf{q^{in}_{1}},\ldots,\mathsf{q^{in}_{m}}\}$ denote a set of queries with trigger sequence $\mathsf{s_{trg}}$ present in each query. We define $\mathrm{P_{top}}$ as the set of top-$k$ passages retrieved in response to a user query $\mathsf{q^{in}_{j}}\in\mathrm{Q_{in}}$, with $\mathsf{p_{adv}}$ included within $\mathrm{P_{top}}$. Let $\mathrm{T_{in}}=\{\mathsf{t_{1:w_{1}}},\ldots,\mathsf{t_{1:w_{m}}}\}$ denote the tokenized input, where $\mathsf{t_{1:w_{j}}}$ represents the tokenized version of the user query $\mathsf{q^{in}_{j}}$ and its passages $\mathrm{P_{top}}$ structured in the pre-specified RAG template format. We use $\mathsf{t_{gen}}\subset\mathsf{t_{1:w_{j}}}$ to denote the subset of tokens that correspond to the string $\mathsf{s_{gen}}$ and $\mathsf{t_{w_{j}+1:w_{j}+c}}$ to represent the tokenized version of the target string $\mathsf{s_{op}}$. We can now formulate the loss function as
(2) $\mathsf{L_{gen}}(\mathrm{T_{in}})=-\sum_{j=1}^{\mathsf{m}}\log\Pr(\mathsf{t_{w_{j}+1:w_{j}+c}}|\mathsf{t_{1:w_{j}}})=-\sum_{j=1}^{\mathsf{m}}\sum_{i=1}^{c}\log\Pr(\mathsf{t_{w_{j}+i}}|\mathsf{t_{1:w_{j}+i-1}})$
@@
where $\Pr(\mathsf{t_{w_{j}+i}}|\mathsf{t_{1:w_{j}+i-1}})$ denotes the probability of generating token $\mathsf{t_{w_{j}+i}}{}$ given the sequence of tokens up to position $\mathsf{w_{j}+i-1}$. We now formulate an optimization problem to solve Equation 2.
ii) Optimization Algorithm.
@@
One potential strategy could be to implement GCG optimization (Zou et al., 2023) to minimize our loss. However, we observe that GCG requires a large batch size and many iterations to converge and break the generator’s alignment. Instead, we propose a more efficient approach, called Multi Coordinate Gradient (MCG), that reduces the number of iterations and batch size. Algorithm 1 (Section 6.1.1) presents our method, which simultaneously updates a subset of $C$ coordinates per iteration. We generate a batch of $B$ candidates, and for each of them we update $C$ random coordinates with one of the top-$k$ token replacements. The candidate with the lowest $\mathsf{L_{gen}}$ is then chosen as the best substitution for $\mathsf{t_{gen}}$.
@@
We observe that loss fluctuates across iterations when C is large. Therefore, after each iteration, we progressively halve the number of coordinates to steadily decrease the $\mathsf{L_{gen}}$. While a large batch size $B$ (e.g., 512) and many iterations (e.g., 128) can also help decrease the loss, this approach is slower and conflicts with our efficiency goal. For success in fewer iterations (e.g., 4 to 16), changing at least 2 coordinates per iteration ($\mathsf{c_{min}}=2$ in Step 10) helps mitigate loss fluctuations while maintaining a manageable batch size of 64. Note that the GCG attack is a special case of our MCG approach when the number of coordinates is set to $C=1$. In Section 6.1 (Table 7), we empirically show that our approach achieves higher attack success than GCG, especially when the adversary has a small iteration budget of $I=4$ and batch size of $B=64$. Table 1 provides examples of adversarial strings $\mathsf{s_{cmd}}$ and $\mathsf{s_{op}}$ that are used to achieve various objectives given in Section 3.1.
@@
Table 1. Concrete Instantiation of Phantom Framework: Examples of adversarial strings $\mathsf{s_{cmd}}$ and $\mathsf{s_{op}}$ used to achieve different adversarial objectives.
@@ 5. Evaluation
All experiments are repeated three times with different random seeds, and all reported values are the averages of these results. Experiments were run on machines with at least 256GB of memory available and different GPU configurations: Nvidia RTX 4090, Nvidia RTX A6000, and Nvidia A100.
Datasets. To evaluate our process, we use the MS MARCO question and answer dataset (Nguyen et al., 2016), which contains more than 8 million document passages and roughly 1 million real user queries. For completeness, we also test our attack on two other datasets, namely Natural Question (NQ) (Kwiatkowski et al., 2019) and HotPot-QA (Yang et al., 2018).
Retrievers and Generators. We evaluate on a variety of commonly used, open-source, retriever and generator models. In particular, we use Contriever (Gautier et al., 2022), Contriever-MS (a version of Contriever fine-tuned on the MS MARCO dataset), and DPR (Karpukhin et al., 2020) as retrievers. For generators, we use the following LLMs: Vicuña (Chiang et al., 2023) version 1.5 7B and 13B, Llama 3 Instruct 8B (Touvron et al., 2023; Meta, 2024), and Gemma Instruct 2B and 7B (Team et al., 2024). Section 6.1.10 presents experiments on larger closed-source models.
@@
RAG Hyperparameters. For our main experiment we set the number of passages retrieved to a default value $k=5$. Our test are conducted on 25 queries selected from the test set, which ensures that the trigger appears naturally within each query. Sections 6.1.6 and 6.1.9 present ablation studies where we vary these parameters. We test for three triggers, for our adversarial objective, where we choose common brand or celebrity names as they frequently appear in natural user queries, and our attack can cause significant reputation damage.
Table 2. Effectiveness of Phantom Attack for Refusal to Answer and Biased Opinion objectives: Attack success reported for two objectives, over three triggers, in scenarios where the adversarial document is retrieved versus No-attack conditions. The RAG uses Contriever for retrieval and one of four LLMs as the generator. Symbol denotes the RAG breaks alignment with only the adversarial command $\mathsf{s_{cmd}}$, while indicates also the need of MCG optimization to break RAG’s alignment.
5.1. Phantom attack on RAG System
Here we evaluate the performance of Phantom on our adversarial objectives. We assess how different RAG generator and retriever hyper-parameters affect our attack’s performance in Section 6.
@@ 5.1. Phantom attack on RAG System
Table 2 reports the results for our Refusal to Answer and Biased Opinion objectives. We observe that all four generators execute the command with high success rate, without requiring to run the MCG optimization. We believe the adversary’s ‘Refusal to Answer’ objective aligns with RAG’s tendency to refrain from answering when uncertain about the user’s prompt, making it easier to succeed.
@@
For the Biased Opinion objective, we are able to induce Gemma-2B and Llama3-8B to produce a biased opinion, with only the adversarial command $\mathsf{s_{cmd}}$ and no MCG optimization. However, for Vicuna-7B and Gemma-7B, the adversarial command $\mathsf{s_{cmd}}$ alone is not sufficient to consistently break the alignment. Consequently, we run 16 iterations of MCG optimization over 16 tokens to successfully break the alignment of these models as well. We observe a significant improvement of $38.7\%$ and $38.4\%$ in the attack’s success across the three triggers when we append MCG’s $\mathsf{s_{gen}}$ to the adversarial command $\mathsf{s_{cmd}}$ for Vicuna-7B and Gemma-7B, respectively. A detailed comparison is given in Table 4 (Section 6.1), which shows the attack success with and without MCG optimization. We include an alternative version of Table 2 in Appendix A.1, showing the average number of queries for which the attack was successful (with related standard deviations) across three runs. Concrete examples for this objective, can be found in Appendix A.4.1.